DDS-BASED FAST SCAN SPECTROMETER

E. A. ALEKSEEV, Institute of Radio Astronomy of NASU, Chervonopraporna 4, 61002 Kharkov, Ukraine; \underline{R. A. MOTIYENKO}, L. MARGULÈS, Laboratoire PhLAM, CNRS UMR 8523, Université de Lille 1, 59655 Villeneuve d’Ascq Cedex, France.

The technique of direct digital synthesis (DDS) has two important features which enable its application in microwave spectroscopy: micro-Hz tuning resolution and extremely fast frequency switching with continuous phase. We have applied a direct digital synthesizer in a PLL-spectrometer based on backward-wave oscillator (BWO). As result we have obtained an instrument that can cover a 100 GHz bandwidth in less than one hour with high spectral resolution and high precision of frequency measurement. The application of the spectrometer to sub-millimeter wave survey spectra records of several isotopic species of astrophysical molecules (methanol, formamide, methyl formate, aziridine) will be discussed. The support of Université de Lille 1 and le Programme National de Physique Chimie du Milieu Interstellaire is gratefully acknowledged.